OpenStudy (anonymous):

Paul and Erin are playing a board game. The spinner has 4 colors on it. Paul needs to spin red two times in order to win. What is the probability that he spins red two times in a row? 1/12 1/8 2/8 1/16

OpenStudy (anonymous):

assuming all colors are equally likely it would be \[\frac{1}{4}\times \frac{1}{4}\]

OpenStudy (anonymous):

Hey! I remember you!

OpenStudy (anonymous):

i can see you bending over the stove...

OpenStudy (anonymous):

so the answser is 1/16 @satellite73

OpenStudy (anonymous):

yes

OpenStudy (holly00d1248):

Its C because hes spinning twice 1/4*2